OpenStudy (shadowcain):

The area of the trapezoid is 39 ft ^2. What is the height of the trapezoid? ____ ft

OpenStudy (anonymous):

where is the given?

OpenStudy (anonymous):

\[Area of trapezoid = \frac{ Base1 + Base2 }{ 2 } \times height\] So \[39^{2} = \frac{ 11+15 }{ 2 } * height\]

OpenStudy (anonymous):

So Solve for height to get that height is equal to 3

jhonyy9 (jhonyy9):

39 why is squared there ?

jhonyy9 (jhonyy9):

so there 39 not is right squared - ok ?

OpenStudy (shadowcain):

ok, but it don't change the question

OpenStudy (shadowcain):

thanks for the help @m_akram

OpenStudy (anonymous):

you are welcome @ShadowCain